# Geometrically Thick Disks around Kerr Black Holes in a Swirling Universe

### v1

@@ 3 Geometrically Thick Disks
In axisymmetric and stationary spacetimes geometrically thick disks can be modeled as a body of fluids, where the fluid particles are moving on circular orbits around a central compact object. The stress-energy tensor of the disk, $T^{\mu\nu}$, can be written as a sum of the different contributing terms, i.e. the fluid, viscosity, radiation, and magnetic terms. In general, the degree of complexity of the model can increase significantly if the various terms are taken into account. In the Polish doughnut model, the viscosity, radiation, and magnetic terms are neglected, and only the fluid part is taken into account. The accretion disk is therefore modeled as a perfect fluid and due to its simplicity, it offers a suitable model to investigate the qualitative properties of geometrically thick accretion disks. The stress-energy tensor in the Polish doughnut model is given by
@@
where $\rho$ is the rest-mass density, $h$ is the specific enthalpy, $u^{\mu}=(u^{t},0,0,u^{\varphi})$ is the four-velocity of the fluid particles, $p$ is the pressure and $g^{\mu\nu}$ is the contra-variant metric tensor. To construct disk solutions the fundamental conservation laws, i.e. the conservation of energy-momentum and the continuity equation need to be solved,
@@
The continuity equation (9) is always trivially satisfied, since $u^{r}=u^{\theta}=0$ for circular motion and $\partial_{t}=\partial_{\varphi}=0$ due to the axisymmetry and stationarity of the spacetime. By applying the orthogonal projection tensor $h^{\alpha}_{\beta}=\delta^{\alpha}_{\beta}+u^{\alpha}u_{\beta}$ in eq. (8) and rewriting it in terms of the specific angular momentum $\ell$ and the angular velocity $\Omega$, the relativistic Euler equations can be derived,
@@
where $-u_{t}$ is the mass-normalized energy. As a necessary integrability condition the von Zeipel theorem must be assumed, which states that the angular velocity can be written as a function of the specific angular momentum, $\Omega=\Omega(\ell)$. By assuming a polytropic equation of state, where the pressure is solely a function of the density, such as $p=K\rho^{\Gamma}$, with $K,\ \Gamma$ constant, the von Zeipel theorem is satisfied. Hence, the eq. (10) becomes integrable and can be written as
[…]
where the subscript $in$ refers to the inner edge of the accretion disk. Assuming a uniform specific angular momentum distribution of the disk, the integral term regarding the specific angular momentum vanishes and the right-hand side can be defined as an effective potential, $\mathcal{W}=\ln|u_{t}|,\ \mathcal{W}_{in}=\ln|(u_{t})|_{in}$, which corresponds to the combined gravitational and centrifugal potential. The effective potential at the inner edge of the disk, $\mathcal{W}_{in}$, is taken as a free parameter. In order to represent all physically feasible solutions, it will be set to $\mathcal{W}_{in}=0$ in all further calculations. Rewriting the specific enthalpy in eq. (11) in terms of the rest-mass density is leading to,
@@
The rest-mass density can therefore be expressed as a function of the effective potential $\mathcal{W}$. As a consequence, the equi-density and equi-pressure surfaces coincide with the equi-potential surfaces. Since $\partial_{\mu}\mathcal{W}=0$ is satisfied for $a_{\mu}=0$, the motion at the local extrema of the effective potential is geodesic. Furthermore, the pressure gradient vanishes at the local extrema of the effective potential. Minima of $\mathcal{W}$ correspond to maxima of the pressure and therefore indicate a disk center (stable geodesic motion of the fluid particles). Maxima of $\mathcal{W}$ corresponds to the minima of the pressure and mark disk cusps (unstable geodesic motion of the fluid particles). The effective potential $\mathcal{W}$ can be expressed through the metric components and the specific angular momentum as,
@@
Due to the odd $\mathcal{Z}2$ spacetime symmetry in the KBHSU solution, the local extrema of the potential are not located in the equatorial plane for $j\neq 0$. In order to determine the locations of the extrema, one needs to solve therefore the system of equations composed of $\partial_{r}\mathcal{W}=0$ and $\partial_{\theta}\mathcal{W}=0$ for the specific angular momentum $\ell$. The resulting distribution of $\ell$, which solves this system of equations, represents therefore the Keplerian specific angular momentum $\ell_{K}$ of the circular geodesics in that spacetime. For a chosen specific angular momentum $\ell_{0}$ of the disk, the location of the disk center is then given by $(r_{center},\theta_{center})=\{(r,\theta):\ell_{K}(r,\theta)=\ell_{0}\wedge\partial_{r}\ell_{K}(r,\theta)>0\}$ and of a disk cusp by $(r_{cusp},\theta_{cusp})=\{(r,\theta):\ell_{K}(r,\theta)=\ell_{0}\wedge\partial_{r}\ell_{K}(r,\theta)<0\}$.
@@ 4 Circular Orbits and Disk solutions
Figure 5: Specific angular momentum $\ell_{K}$ of circular orbits for the different Kerr parameter solutions with varying swirling parameter $j$. The upper row presents $\ell_{K}$ for prograde orbits and the lower row presents $\ell_{K}$ for retrograde orbits. The terms prograde and retrograde refer here to the rotation direction of the KBHSU, which is defined by the Kerr spin parameter $a$. The dotted vertical line in each plot marks the marginally bound orbit $r_{mb}$ and the dashed-dotted vertical line marks the marginally stable orbit $r_{ms}$ of the Kerr solution ($j=0$). Dashed vertical lines mark the outer marginally stable orbits $r_{ms}^{out}$, which appear for the KBSHU solutions ($j\neq 0$). Black curve sections mark unbound orbits.
The specific angular momentum of circular orbits changes significantly for the KBHSU solutions compared to the Kerr solutions. Up to a critical value of the swirling parameter, $j_{c}$, the specific angular momentum develops an outer extremum. For prograde orbits this extremum is a maximum, for retrograde orbits it is a minimum (Fig. 5 upper and lower row). For radial values beyond the outer extremum, the specific angular momentum converges asymptotically to 0 in both cases. The outer extremum corresponds to the emergence of an outer marginally stable orbit for KBHSU solutions. Stable circular orbits are therefore only possible in a specific spacetime region, which is lower bounded by the inner marginally stable orbit and upper bounded by the outer marginally stable orbit. The inner marginally stable orbits of the KBHSU solutions are closely located to the marginally stable orbit of the Kerr solution. A greater deviation of $\ell_{K}$ from the Kerr solution is only noticeable from the inner extremum outwards. For solutions with a swirling parameter greater than the critical swirling parameter, $j>j_{c}$, the extrema vanishes and the specific angular momentum becomes monotonic, all circular orbits are therefore unstable. For prograde orbits this critical value $j_{c}$ increases with increasing Kerr spin parameter $a$, therefore stable circular orbits are possible for a larger range of the swirling parameter $j$ (Fig. 5 upper row). For retrograde orbits the value of $j_{c}$ decreases with increasing Kerr spin parameter $a$, stable circular orbits for fast rotating KBHSU occur therefore for a smaller range of the swirling parameter (Fig. 5 lower row). The swirling rotation causes furthermore the emergence of static orbits for prograde as well as retrograde motion. An illustration of the rest specific angular momentum, which is used to determine the static orbits, can be found in Fig. 6.
@@
Figure 6: Rest specific angular momentum, $\ell_{r}=-\frac{g_{t\varphi}}{g_{tt}}$, plotted alongside the prograde and retrograde circular orbits for the various KBHSU solutions. The previously shown $\ell_{K}$ distributions are plotted as dotted curves for comparison. The circular marker symbols indicate the radial location of the static orbits, $r_{so}$. At these radial locations $\ell_{K}$ is identical to $\ell_{r}$ and as a consequence, the corresponding circular orbit is a static orbit.
The rest specific angular momentum $\ell_{r}$ alongside the circular orbits has an intersection with $\ell_{K}$ for the analyzed KBHSU solutions. The specific angular momentum of the circular orbit corresponding to this intersection is equal to the specific angular momentum a particle at rest would have in the co-moving reference frame. Therefore a test particle, which is initially at rest in the co-moving reference frame, would always stay at rest in this reference frame in the absence of disturbances. However, it should be noted that all static orbits are unstable since the intersection point for each solution is outwards of $r_{ms}^{out}$. Furthermore, for $r>r_{so}$ the absolute value of $\ell_{K}$ is smaller than $\ell_{r}$ for prograde as well as retrograde motion. Thus, all circular orbits outwards of the static orbit are counter-rotating in the co-moving reference frame of the spacetime background. With increasing $j$ the radial location $r_{so}$ of the static orbits is moving inwards, regardless of the Kerr parameter $a$. The differences between the various Kerr parameters are in general negligible, $r_{so}$ is located for each solution roughly at the same radial location for a fixed $j$, regardless of the rotational direction. Therefore we conclude, that the Kerr parameter and rotation direction have a lesser impact on the properties of static orbits, which are mostly determined by $j$.
As mentioned in section 3, the circular geodesics are not located in the equatorial plane, in contrast to the Kerr solution. To each point of the presented curves for the $\ell_{K}$ and $\ell_{r}$ distributions, there is a corresponding varying $\theta$ coordinate. For a better understanding regarding the spatial distribution of the circular orbits, a cross-section plot is depicted in Fig. 7 for the various analyzed KBHSU solutions.
@@
Figure 7: Cross section plots of circular orbits and their corresponding specific angular momentum $\ell_{K}$ for the different Kerr parameter solutions with varying swirling parameter $j$. The upper row presents $\ell_{K}$ of prograde orbits and the lower row presents $\ell_{K}$ for retrograde orbits. The terms prograde and retrograde refer here to the rotation direction of the KBHSU, which is defined by the Kerr spin parameter $a$. The different symbols mark selected solutions for circular orbits for varying $j$. They are coloured according to their specific angular momentum, which is described by the scale on the right of each plot. Unstable orbits are marked by a black edge around the symbol and the static orbits are marked by a circle.
@@
The spatial location of KBHSU circular orbits is highly affected by the swirling parameter, it deviates from the equatorial plane and is located in planes orthogonal to the rotational axis in the upper and lower hemispheres. With increasing $j$ the verticality of the circular orbit distribution increases and for $\ell_{K}\rightarrow 0$ the height of the corresponding circular orbit above or below the equatorial plane diverges. The location of the static orbits is largely unaffected by $a$ and for increasing $j$ it moves closer to the equatorial plane. Stable orbits are only possible close to the equatorial plane. In the case of prograde motion, the orbits are solely located in the upper hemisphere (Fig. 7 upper row), they are therefore co-moving with the swirling background. For retrograde motion, the orbits are solely located in the lower hemisphere (Fig. 7 lower row). Since the frame-dragging direction of the swirling background changes its sign for the lower hemisphere, a counter-rotating test particle regarding the Kerr rotation is co-moving with the swirling background in the lower hemisphere. All circular orbits are therefore co-moving with their surrounding spacetime background reference frame. The spatial distribution of circular orbits throughout the Kerr parameter range is very similar, the influence of the Kerr parameter is negligible. We conclude, that the general properties of circular orbits for KBHSU are largely determined by the swirling parameter, with the Kerr parameter playing a minor role. However, the Kerr parameter has a significant influence on the angular momentum distribution and the stability of the circular orbits. The rotation of the KBHSU has a stabilizing effect on the co-rotating orbits and a destabilizing effect on the counter-rotating orbits. This can be attributed to the rotation of the swirling background. Orbits located in the lower-hemisphere are co-rotating with the spacetime background but counter-rotating with the black hole, which is inducing instabilities due to the opposite spin direction of the swirling and Kerr rotation. The magnitude of this effect grows with increasing swirling and Kerr parameters $j$ and $a$.
@@
As a consequence accretion disks with an uniform constant angular momentum can exist only for a specific range of $\ell_{0}$, which is lower bounded by $\ell_{ms}^{in}$ and upper bounded by $\ell_{ms}^{out}$, since only for $\ell_{0}\in[\ell_{ms}^{in},\ell_{ms}^{out}]$ stable circular orbits exist. Thus, disk solutions with static orbits and static surfaces do not exist. The disk center can only be located between $r_{ms}^{in}$ and $r_{ms}^{out}$ and with increasing $j$ the solution space for prograde accretion tori gets smaller, as the solution range of stable orbits decreases. Due to the stabilizing effect of the KBHSU rotation, the solution space for prograde (retrograde) accretion tori is greater (smaller) for faster rotating KBHSUs compared to slower rotating ones. This is depicted exemplary in Fig. 8 for $j=10^{-4}$ and varying values for $\ell_{0}$.
@@
Figure 8: Effective potential $\mathcal{W}$ alongside circular orbits for the different Kerr spin parameters and various $\ell_{0}$. The swirling parameter is set to $j=10^{-4}$ in all solutions. The upper row presents $\mathcal{W}$ for prograde motion and the lower row for retrograde motion (referring to the rotation direction defined by the Kerr parameter $a$). Dashed curves indicate solutions for which no disks can exist, due to the absence of a local minimum.
With increasing $a$ prograde disk exists for a greater range of $\ell_{0}$ for the same $j$ (Fig. 8 upper row). For retrograde disks, the solution space narrows down for the same $j$ and $\ell_{0}$ parameter range (Fig. 8 lower row). All possible disk solutions can be classified into two main types: Type 1 solutions are composed of an inner cusp and an outer cusp, between which the disk center is located. The inner cusp corresponds to an inner local maximum, the disk center to the local minimum and the outer cusp to the outer local maximum of the effective potential $\mathcal{W}$. These solutions can be further classified into three subtypes: Type 1a disks, where the inner maximum is greater than the outer maximum (e. g. Fig. 8 (d)). Type 1b disks, where the outer maximum is greater than the inner maximum (e. g. Fig. 8 (b)). And the special case Type 1c, where the inner and outer potential maxima have the same value. It can be viewed as the limiting case of a Type 1a or 1b solution. In contrast to Type 1 solutions, Type 2 solutions are only composed of a disk center and an outer cusp, the effective potential has therefore one minimum and only one maximum, located to the outside of the minimum (e. g. Fig. 8 (c)). With increasing $a$ the disk solution space moves from mostly Type 1 solutions to Type 2 solutions for prograde disks. For retrograde disks all solutions are of Type I. Exemplary cross-section plots of the different disk types are presented in Fig. 9 for prograde disks. In the case of retrograde disk, the defined disk types are analogous, besides that only solutions of Type 1 exist.
@@
For Type 1a disks the equi-potential surface of the inner cusp is not closed, it is open towards the KBHSU center and expands from the cusp outwards to infinity. The equi-potential surface of the outer cusp is semi-closed, with the closed region containing the disk center and the open region expanding from the outer cusp towards infinity. This enclosed region with the disk center is fully surrounded by the equi-potential surface of the inner cusp. Depending on the initial conditions various disk evolutions seem possible, as the complexity of disk dynamics could increase due to the presence of two cusps. Matter could be accreted through the inner cusp by flowing past the center region, without having a major effect on the dynamics and stability of the center region, as it is fully enclosed by the equi-potential surface of the outer cusp. On a different note, the whole disk could become unstable, since matter from the disk center could disperse through the outer cusp to unstable orbits, leading to a diffusion of the whole inner disk region. In case of Type 1b disks, the equi-potential surface of the inner cusp is semi-closed, it fully encloses the disk center and is open towards the KBHSU center. The equi-potential surface of the outer cusp is open towards the KBHSU center and from the outer cusp outwards to infinity. It surrounds the inner cusp potential surface and the disk center. Accretion processes through both cusps are conceivable. Matter from outside could accrete through the outer cusp towards the KBHSU center, flowing past the center region. Accretion from the disk center through the inner cusp towards the KBHSU could also take place. Since the center region is fully surrounded by the outer cusp potential surface, an indirect accretion process could take place, where matter flows through the outer cusp towards the inner potential surface, feeding the central region of the accretion disk, which could trigger an accretion process through the inner cusp. In the limiting case Type 1c, the effective potential at the inner and outer cusp is identical, the corresponding equi-potential surface has therefore two intersections and is open towards the KBHSU center and open outwards of the outer cusp. Between the inner and outer cusp it fully encloses the disk center region. An accretion process may happen through the outer cusp towards the disk center and through the inner cusp towards the KBSHU center. The disk could also disperse if matter from the center region flows through the outer cusp to the outer regions. For Type 2 disks only an outer cusp exists, with the equi-potential surface of the cusp being closed towards the KBHSU and open outwards of the cusp towards infinity. As for Type 1a disks it fully encloses the disk center. Due to the absence of an inner cusp, no accretion processes without additional inner-physical disk effects are possible. Nevertheless, the discussed runaway instability for Type 1 disks could also apply to Type 2 disks, where matter dispersing through the outer cusp towards unstable orbits could lead to a diffusion of the center region.
@@
In order to investigate the influence of the Kerr parameter on prograde disk properties, different exemplary solutions for the same parameters, besides the Kerr parameter, are presented in Fig. 10.

### v2

@@ 3 Geometrically Thick Disks
In axisymmetric and stationary spacetimes, geometrically thick disks can be modeled as a body of fluids, where the fluid particles are moving on circular orbits around a central compact object. The stress-energy tensor of the disk, $T^{\mu\nu}$, can be written as a sum of the different contributing terms, i.e. the fluid, viscosity, radiation, and magnetic terms. In general, the degree of complexity of the model can increase significantly if the various terms are taken into account. In the Polish doughnut model, the viscosity, radiation, and magnetic terms are neglected and only the fluid part is taken into account. The accretion disk is therefore modeled as a perfect fluid and due to its simplicity, it offers a suitable model to investigate the qualitative properties of geometrically thick accretion disks. The stress-energy tensor is then given by
@@
where $\rho$ is the rest mass density, $h$ is the specific enthalpy, $u^{\mu}=(u^{t},0,0,u^{\varphi})$ is the four-velocity of the fluid particles, $p$ is the pressure and $g^{\mu\nu}$ is the contra-variant metric tensor. To construct disk solutions, the fundamental conservation laws, i.e. the conservation of energy-momentum and the continuity equation need to be solved,
@@
The continuity equation (9) is always trivially satisfied since $u^{r}=u^{\theta}=0$ for circular motion and $\partial_{t}=\partial_{\varphi}=0$ due to the axisymmetry and stationarity of the spacetime. By applying the orthogonal projection tensor $h^{\alpha}_{\beta}=\delta^{\alpha}_{\beta}+u^{\alpha}u_{\beta}$ to eq. (8) and by rewriting it in terms of the specific angular momentum $\ell=-\frac{u_{\varphi}}{u_{t}}$ and the angular velocity $\Omega=\frac{u^{\varphi}}{u^{t}}$ of the particle, the relativistic Euler equations can be derived,
@@
where $-u_{t}$ is the mass-normalized energy. As a necessary integrability condition the von Zeipel theorem must be presumed, which states that the angular velocity can be written as a function of the specific angular momentum, $\Omega=\Omega(\ell)$. By assuming a polytropic equation of state, where the pressure is solely a function of the density, such as $p=K\rho^{\Gamma}$, with $K$ and $\Gamma$ as constants, the von Zeipel theorem is always satisfied. Hence, the eq. (10) becomes integrable and can be written as
$\ln|u_{t}|-\ln|(u_{t})_{in}|-\int_{\ell_{in}}^{\ell}\frac{\Omega}{1-\Omega\ell^{\prime}}{\rm d}\ell^{\prime}=-\int_{p_{in}}^{p}\frac{{\rm d}p^{\prime}}{\rho h}=-\int_{h_{in}}^{h}\frac{{\rm d}h^{\prime}}{h^{\prime}}=-\ln h,$ (11)
where the subscript $in$ refers to the inner edge of the accretion disk. Assuming a uniform specific angular momentum distribution of the disk, the integral term regarding the specific angular momentum vanishes and the right-hand side can be defined as an effective potential, $\mathcal{W}=\ln|u_{t}|$ and $\mathcal{W}_{in}=\ln|(u_{t})|_{in}$, which corresponds to the combined gravitational and centrifugal potential of a disk particle. The effective potential at the inner edge of the disk $\mathcal{W}_{in}$ is taken as a free parameter which should not be lower than $\mathcal{W}_{center}$ and not higher than 0, since particles with $\mathcal{W}>0$ could escape to infinity. In order to represent all physically feasible solutions (composed of particles on bound orbits), we will set $\mathcal{W}_{in}=0$ in all further calculations. Rewriting the specific enthalpy in eq. (11) in terms of the rest mass density is leading to,
@@
The rest mass density can therefore be expressed as a function of the effective potential $\mathcal{W}$. As a consequence, the equi-density and equi-pressure surfaces coincide with the equi-potential surfaces. The four-acceleration $a_{\mu}$ can be expressed through the partial derivatives of the effective potential, $a_{\mu}=\partial_{\mu}\mathcal{W}$. Thus, the extrema of the effective potential correspond to geodesic motion. Minima of $\mathcal{W}$ correspond to stable geodesic motion, where the pressure and rest mass density are maximal. The minima are defined as the accretion disk center. Maxima of $\mathcal{W}$ correspond to unstable geodesic motion and correspond to minima of the pressure and rest mass density, they mark the disk cusps. Cusps are characterized by a self-intersection of the equi-potential surface. Due to the unstable flow at their radial location, small perturbations could trigger accretion processes. By expressing $u_{t}$ through the metric components, the effective potential $\mathcal{W}$ is solely determined by the spacetime geometry and the specific angular momentum of the disk particles,
@@
Due to the odd $\mathcal{Z}_{2}$ spacetime symmetry in the KBHSB solution, the local extrema of the potential are not located in the equatorial plane for $j\neq 0$. In order to determine the locations of the extrema, one needs to solve the system of equations composed of $\partial_{r}\mathcal{W}=0$ and $\partial_{\theta}\mathcal{W}=0$ for the specific angular momentum $\ell$. The resulting distribution of $\ell$, which solves this system of equations, represents therefore the Keplerian specific angular momentum $\ell_{K}$ of the circular geodesics for that spacetime. For a chosen specific angular momentum $\ell_{0}$ of the disk, the location of the disk center is then given by $(r_{center},\theta_{center})=\{(r,\theta):\ell_{K}(r,\theta)=\ell_{0}\wedge\partial_{r}|\ell_{K}(r,\theta)|>0\}$ and of a disk cusp by $(r_{cusp},\theta_{cusp})=\{(r,\theta):\ell_{K}(r,\theta)=\ell_{0}\wedge\partial_{r}|\ell_{K}(r,\theta)|<0\}$.
@@ 4 Circular Orbits and Disk solutions
Figure 5: Specific angular momentum $\ell_{K}$ of circular orbits for the different Kerr parameter solutions with varying swirling parameter $j$. The upper row presents $\ell_{K}$ for prograde orbits, and the lower row presents $\ell_{K}$ for retrograde orbits. The terms prograde and retrograde refer here to the rotation direction of the KBHSB, which is defined by the Kerr spin parameter $a$. The dotted vertical line in each plot marks the marginally bound orbit $r_{mb}$ and the dashed-dotted vertical line marks the marginally stable orbit $r_{ms}$ of the Kerr solution ($j=0$). Dashed vertical lines mark the outer marginally stable orbits $r_{ms}^{out}$, which appear for the KBHSB solutions ($j\neq 0$). Black curve sections mark unbound orbits.
The specific angular momentum of circular orbits changes significantly for the KBHSB solutions compared to the Kerr solutions. Up to a critical value of the swirling parameter, $j_{c}$, the specific angular momentum develops an outer extremum. For prograde orbits, this extremum is a maximum; for retrograde orbits, it is a minimum (Fig. 5 upper and lower row, respectively). For radial values beyond the outer extremum, the specific angular momentum converges asymptotically to 0 in both cases. The outer extremum corresponds to the emergence of an outer marginally stable orbit for KBHSB solutions. Stable circular orbits are therefore only possible in a specific spacetime region, which is lower bounded by the inner marginally stable orbit and upper bounded by the outer marginally stable orbit. The inner marginally stable orbits of the KBHSB solutions are closely located to the marginally stable orbit of the Kerr solution. A greater deviation of $\ell_{K}$ from the Kerr solution is only noticeable from the inner extremum outwards. For solutions with a swirling parameter greater than the critical swirling parameter, $j>j_{c}$, the extrema vanish and the specific angular momentum becomes monotonic; all circular orbits are therefore unstable. For prograde orbits, the critical value $j_{c}$ increases with increasing Kerr spin parameter $a$, thus stable circular orbits are possible for a larger range of the swirling parameter $j$ (Fig. 5 upper row). For retrograde orbits, the value of $j_{c}$ decreases with increasing Kerr spin parameter $a$; stable circular orbits for fast rotating KBHSB occur therefore for a smaller range of the swirling parameter (Fig. 5 lower row). Furthermore, the swirling rotation causes the emergence of static orbits for prograde as well as retrograde motion. An illustration of the rest specific angular momentum, which is used to determine the static orbits can be found in Fig. 6.
@@
Figure 6: Rest specific angular momentum, $\ell_{r}=-\frac{g_{t\varphi}}{g_{tt}}$, plotted alongside the prograde and retrograde circular orbits for the various KBHSB solutions. The previously shown $\ell_{K}$ distributions are plotted as dotted curves for comparison. The circular marker symbols indicate the radial location of the static orbits $r_{so}$. At these radial locations $\ell_{K}$ is identical to $\ell_{r}$, as a consequence the corresponding circular orbit is a static orbit.
The rest specific angular momentum $\ell_{r}$ alongside the circular orbits has an intersection with $\ell_{K}$ for the analyzed KBHSB solutions. The specific angular momentum of the circular orbit corresponding to this intersection is equal to the specific angular momentum a particle at rest would have in the co-moving reference frame. Therefore, a test particle, which is initially at rest in the co-moving reference frame, would always stay at rest in this reference frame in the absence of disturbances. However, it should be noted that all static orbits are unstable since the intersection point for each solution is outwards of $r_{ms}^{out}$. Furthermore, for $r>r_{so}$, the absolute value of $\ell_{K}$ is smaller than $\ell_{r}$ for prograde as well as retrograde motion. Thus, all circular orbits outwards of the static orbit are counter-rotating in the co-moving reference frame of the spacetime background. With increasing $j$, the radial location $r_{so}$ of the static orbits is moving inwards, regardless of the Kerr parameter $a$. The differences between the various Kerr parameters are in general negligible, thus $r_{so}$ is located for each solution roughly at the same radial location for a fixed $j$, regardless of the rotational direction. Therefore, we conclude that the Kerr parameter and rotation direction have a lesser impact on the properties of static orbits, which are mostly determined by $j$.
As mentioned in section 3 the circular geodesics are not located in the equatorial plane, in contrast to the Kerr solution. To each point of the presented curves for the $\ell_{K}$ and $\ell_{r}$ distributions, there is a corresponding varying $\theta$ coordinate. For a better understanding regarding the spatial distribution of the circular orbits, a cross-section plot is shown in Fig. 7 for the various analyzed KBHSB solutions.
@@
Figure 7: Cross section plots of circular orbits and their corresponding specific angular momentum $\ell_{K}$ for the different Kerr parameter solutions with varying swirling parameter $j$. The upper row presents $\ell_{K}$ of prograde orbits and the lower row presents $\ell_{K}$ for retrograde orbits. The terms prograde and retrograde refer here to the rotation direction of the KBHSB, which is defined by the Kerr spin parameter $a$. The different symbols mark selected solutions for circular orbits for varying $j$. They are coloured according to their specific angular momentum, which is described by the scale on the right of each plot. Unstable orbits are marked by a black edge around the symbol, and the static orbits are marked by a circle.
@@
The spatial location of KBHSB circular orbits is highly affected by the swirling parameter, it deviates from the equatorial plane and is located in planes orthogonal to the rotational axis in the upper and lower hemispheres. With increasing $j$ the verticality of the circular orbit distribution increases, and for $\ell_{K}\rightarrow 0$ the height of the corresponding circular orbit above or below the equatorial plane diverges. The location of static orbits is largely unaffected by $a$ and for increasing $j$ it moves closer to the equatorial plane. Stable orbits are only possible close to the equatorial plane. In the case of prograde motion, the orbits are solely located in the upper hemisphere (Fig. 7 upper row); they are therefore co-moving with the swirling background. For retrograde motion, the orbits are solely located in the lower hemisphere (Fig. 7 lower row). Since the frame-dragging direction of the swirling background changes its sign for the lower hemisphere, a counter-rotating test particle regarding the Kerr rotation is co-moving with the swirling background in the lower hemisphere. All circular orbits are therefore co-moving with their surrounding spacetime background reference frame. The spatial distribution of circular orbits throughout the Kerr parameter range is very similar, thus the influence of the Kerr parameter is negligible. We conclude that the general properties of circular orbits for KBHSB are largely determined by the swirling parameter, with the Kerr parameter playing a minor role. However, the Kerr parameter has a significant influence on the values of the angular momentum distribution and the stability of the circular orbits. The rotation of the KBHSB has a stabilizing effect on the co-rotating orbits and a destabilizing effect on the counter-rotating orbits. This can be attributed to the rotation of the swirling background. Orbits located in the lower hemisphere are co-rotating with the spacetime background but counter-rotating with the black hole, which is inducing instabilities due to the opposite spin direction of the swirling and Kerr rotation. The magnitude of this effect grows with increasing swirling and Kerr parameters $j$ and $a$.
@@
As a consequence, accretion structures can exist only for a specific range of $\ell_{0}$, which is lower bounded by $\ell_{ms}^{in}$ and upper bounded by $\ell_{ms}^{out}$, since only for $|\ell_{0}|\in[|\ell_{ms}^{in}|,|\ell_{ms}^{out}|]$ stable circular orbits exist. Thus, disk solutions with static orbits and static surfaces do not exist. The disk center can only be located between $r_{ms}^{in}$ and $r_{ms}^{out}$ and with increasing $j$ the solution space for prograde accretion structures gets smaller, as the solution range of stable orbits decreases. Due to the stabilizing effect of the KBHSB rotation, the solution space for prograde (retrograde) accretion structures is greater (smaller) for faster rotating KBHSBs compared to slower rotating ones. This is depicted exemplary in Fig. 8 for $j=10^{-4}$ and varying values for $\ell_{0}$.
@@
Figure 8: Effective potential $\mathcal{W}$ alongside circular geodesics for the different Kerr spin parameters and various $\ell_{0}$. The swirling parameter is set to $j=10^{-4}$ in all solutions. The upper row presents $\mathcal{W}$ for prograde motion and the lower row for retrograde motion (referring to the rotation direction defined by the Kerr parameter $a$). Dashed curves indicate solutions for which no closed disk surfaces can exist due to the absence of a local minimum.
With increasing $a$ prograde disk exists for a greater range of $\ell_{0}$ for the same $j$ (Fig. 8 upper row). For retrograde disks the solution space narrows down for the same $j$ and $\ell_{0}$ parameter range (Fig. 8 lower row). All possible disk solutions can be classified into two main types: Type 1 solutions are composed of an inner cusp and an outer cusp, between which the disk center is located. The inner cusp corresponds to an inner local maximum, the disk center to the local minimum, and the outer cusp to the outer local maximum of the effective potential $\mathcal{W}$. These solutions can be further classified into three subtypes: Type 1a disks, where the inner maximum is greater than the outer maximum (Fig. 8 (d)). Type 1b disks, where the outer maximum is greater than the inner maximum (Fig. 8 (b)). And the special case Type 1c, where the inner and outer potential maxima have the same value. It can be viewed as the limiting case of a Type 1a or 1b solution. In contrast to Type 1 solutions, Type 2 solutions are only composed of a disk center and an outer cusp, the effective potential has one minimum and only one maximum, which is located outwards to the minimum (Fig. 8 (c)). With increasing $a$ the disk solution space moves from mostly Type 1 solutions to Type 2 solutions for prograde disks. For retrograde disks all solutions are of Type I. Exemplary cross-section plots of the different disk types are presented in Fig. 9 for prograde disks. In the case of retrograde disks, the defined disk types are analogous; besides that only solutions of Type 1 exist as mentioned.
@@
For Type 1a disks, the equi-potential surface of the inner cusp is not closed, it is open towards the KBHSB center and expands from the cusp outwards to infinity. The equi-potential surface of the outer cusp is semi-closed, with the closed region containing the disk center and the open region expanding from the outer cusp towards infinity. This enclosed region with the disk center is fully surrounded by the equi-potential surface of the inner cusp. Depending on the initial conditions various disk evolutions seem possible, as the complexity of disk dynamics could increase due to the presence of two cusps. Matter could be accreted through the inner cusp by flowing past the center region, without having a major effect on the dynamics and stability of the center region, as it is fully enclosed by the equi-potential surface of the outer cusp. On a different note, the whole disk could become unstable, since matter from the disk center could disperse through the outer cusp to unstable orbits, leading to a diffusion of the whole inner disk region. In the case of Type 1b disks, the equi-potential surface of the inner cusp is semi-closed, it fully encloses the disk center and is open towards the KBHSB center. The equi-potential surface of the outer cusp is open towards the KBHSB center and from the outer cusp outwards to infinity. It surrounds the inner cusp potential surface and the disk center. Accretion processes through both cusps are conceivable. Matter from outside could accrete through the outer cusp towards the KBHSB center, flowing past the center region. Accretion from the disk center through the inner cusp towards the KBHSB could also take place. Since the center region is fully surrounded by the outer cusp potential surface, an indirect accretion process could take place, where matter flows through the outer cusp towards the inner potential surface, feeding the central region of the accretion structure, which could trigger an accretion process through the inner cusp. In the limiting case Type 1c, the effective potential at the inner and outer cusp is identical; the corresponding equi-potential surface has two intersections, it is open towards the KBHSB center and open outwards of the outer cusp. Between the inner and outer cusp it fully encloses the disk center region. An accretion process may happen through the outer cusp towards the disk center and through the inner cusp towards the KBHSB center. The disk could also disperse if matter from the center region flows through the outer cusp to the outer regions. For Type 2 disks only an outer cusp exists, with the equi-potential surface of the cusp being closed towards the KBHSB and open outwards of the cusp towards infinity. As for Type 1a disks it fully encloses the disk center. Due to the absence of an inner cusp, no accretion processes without additional inner-physical disk effects are possible. Nevertheless, the discussed runaway instability for Type 1 disks could also apply to Type 2 disks, where matter dispersing through the outer cusp towards unstable orbits could lead to a diffusion of the center region.
@@
In order to investigate the influence of the Kerr parameter on prograde disk properties, different exemplary solutions for the same parameters and varying Kerr parameter are presented in Fig. 10.